# Consistency Policy Accelerated Visuomotor Policies via Consistency Distillation

### v1

@@ Abstract
Many robotic systems, such as mobile manipulators or quadrotors, cannot be equipped with high-end GPUs due to space, weight, and power constraints. These constraints prevent these systems from leveraging recent developments in visuomotor policy architectures that require high-end GPUs to achieve fast policy inference. In this paper, we propose Consistency Policy, a faster and similarly powerful alternative to Diffusion Policy for learning visuomotor robot control. By virtue of its fast inference speed, Consistency Policy can enable low latency decision making in resource-constrained robotic setups. A Consistency Policy is distilled from a pretrained Diffusion Policy by enforcing self-consistency along the Diffusion Policy’s learned trajectories. We compare Consistency Policy with Diffusion Policy and other related speed-up methods across 6 simulation tasks as well as two real-world tasks where we demonstrate inference on a laptop GPU. For all these tasks, Consistency Policy speeds up inference by an order of magnitude compared to the fastest alternative method and maintains competitive success rates. We also show that the Conistency Policy training procedure is robust to the pretrained Diffusion Policy’s quality, a useful result that helps practioners avoid extensive testing of the pretrained model. Key design decisions that enabled this performance are the choice of consistency objective, reduced initial sample variance, and the choice of preset chaining steps. Code and training details will be released publicly.
@@ IV-B Simulation Experiments
Table III: Simulation Inference Speeds – Simulation inference speeds were measured on an NVIDIA P5000 datacenter GPU and averaged over 50 rollouts. Benchmarking was done with vanilla Diffusion Policy.
@@ IV-C Real World Experiments
Tasks: We evaluate Consistency Policy in the real world on two tasks: Trash Clean Up and Plug Insertion.
@@
We use a laptop containing a single 3070-Ti GPU with 8GB of VRAM for inference across all real world experiments. Note that we are unable to run ParaDiGMSeffectively on this setup due to ParaDiGMS’s high VRAM usage. We also note that, on the 3070-Ti GPU, the 100-step DDPM variant of Diffusion Policy has an inference time of around $1.5$ seconds per forward pass. Thus, we choose as our baseline method the faster and more realistic DDiM variant of Diffusion Policy, which uses 15-steps for policy inference.
For all methods, we maintain the same policy inputs, outputs, and architecture as in the simulation experiment setup, except the observation size of the plug insertion task: for this task, we use image size of $256$x256256256256. At a given timestep $t$, our policy receives: an over-the-shoulder image, a camera-in-hand image, current end-effector pose and current gripper width. Our policy outputs an action sequence of length $16$, where each step is a $10$-D vector of goal end effector position ($3$-D), goal effector rotation ($6$-D) and goal gripper position ($1$-D).
Metrics: For each trained policy, we report average success rates on one policy checkpoint (selected using validation mean-squared error) and inference time on the 3070-Ti GPU. For success rates, average over 10 trials where we randomize the position of the trash to be picked up.
Results: Table IV shows how the baseline DDiM-variant of Diffusion Policy achieves similar average success rates as our method on the Rubbish Clean Up and Plug Insertion task. However, our method has much lower inference time ($\sim 9$x lower latency).
As Consistency Policy requires 15x fewer forward passes than DDiM, one might expect the full inference speed-up to also be around $15$x instead of around 9x. This discrepancy exists because of overhead costs such as from running the observation encoder. We discuss inference speeds further in the Supplementary Material.
@@
Trash Clean Up Plug Insertion
Success Inference Success Inference
@@
DDiM 0.8 ±.$\pm.13$ 192 $0.6\pm 0.11$ 198
CP (ours) 0.8 ±.$\pm.13$ 21 $0.7\pm 0.10$ 22
Table IV: Real World Experiment Results for Multiple Tasks – Results presented for Consistency Policy and the DDiM variant of Diffusion Policy. Success rates, standard errors and inference speeds are provided for each task, averaged over 10 trials with randomized conditions per trial.
@@ VI Conclusion
We present the use of consistency-based training objectives for training high performing, low-latency visuomotor robot policies. In our evaluation suite of 8 tasks in simulation and in the real world, we demonstrate that consistency policies yield a dramatic increase in inference speeds compared to previous diffusion policy methods without sacrificing success rates. Through our ablations, we also highlight key design decisions that led to visuomotor policies with the highest success rates. These include: the choice of consistency objective, lowering the initial sample variance, the use of dropout, and the choice of preset chaining steps. In future work, we hope to explore deeper explanations behind certain experimental results, such as chaining dynamics and the role of dropout along the $s\rightarrow 0$ generations, and also port Consistency Policy to robots that can better utilize fast inference speeds, including legged, winged, and other mobile platforms.
@@ Appendix B Low Variance Initial Samples
Testing EDM on unconditional image generation suggests that this low-variance sample might have some specific benefit for the robotics domain. Figures 8 and 8 are generations from the same EDM model (trained on CIFAR-10, a $32$x32323232 px image dataset) using low initial variance sampling and high initial variance sampling.
@@ Appendix C Real World Experiment Setup
We perform real world experiments using a Franka Panda robot. We mount a wrist (Zed Mini) camera and an over-the-should camera (Zed Mini) to obtain two images at each timestep. We use an Meta Quest 2 VR setup to collect 180 demonstrations for our task. The Franka robot accepts control commands at 1kHz, and receives new commands from either the VR controller or trained policy at 15 Hz. Note that Consistency Policy generates an end effector action sequence output in around 21ms, and we supply each waypoint to the robot at 15Hz. When the policy performs inference, we send a control command to the robot to maintain its current end effector pose.

### v2

@@ Abstract
Many robotic systems, such as mobile manipulators or quadrotors, cannot be equipped with high-end GPUs due to space, weight, and power constraints. These constraints prevent these systems from leveraging recent developments in visuomotor policy architectures that require high-end GPUs to achieve fast policy inference. In this paper, we propose Consistency Policy, a faster and similarly powerful alternative to Diffusion Policy for learning visuomotor robot control. By virtue of its fast inference speed, Consistency Policy can enable low latency decision making in resource-constrained robotic setups. A Consistency Policy is distilled from a pretrained Diffusion Policy by enforcing self-consistency along the Diffusion Policy’s learned trajectories. We compare Consistency Policy with Diffusion Policy and other related speed-up methods across 6 simulation tasks as well as three real-world tasks where we demonstrate inference on a laptop GPU. For all these tasks, Consistency Policy speeds up inference by an order of magnitude compared to the fastest alternative method and maintains competitive success rates. We also show that the Conistency Policy training procedure is robust to the pretrained Diffusion Policy’s quality, a useful result that helps practioners avoid extensive testing of the pretrained model. Key design decisions that enabled this performance are the choice of consistency objective, reduced initial sample variance, and the choice of preset chaining steps.
@@ IV-B Simulation Experiments
Table III: Simulation Inference Speeds – Simulation inference speeds were measured on an NVIDIA P5000 datacenter GPU and averaged over 50 rollouts. Benchmarking was done with vanilla Diffusion Policy since we used this as our baseline.
@@ IV-C Real World Experiments
Tasks: We evaluate Consistency Policy in the real world on three tasks: Trash Clean Up, Plug Insertion, and Microwave.
@@
  3. 3.
    Microwave: The robot (a mobile manipulator) has to open a microwave, navigate to and retrieve a bag of broccoli, place it in the microwave, close the microwave door, and press the “vegetable” button to start the microwave. This task tests long horizon performance and control of a mobile base along with a standard robot arm (see Fig. 7).
@@
For the first two tasks, we use a laptop containing a single 3070 Ti GPU with 8GB of VRAM for inference. Note that we are unable to run ParaDiGMS effectively on this setup due to ParaDiGMS’s high VRAM usage. We also note that, on the 3070 Ti GPU, the 100-step DDPM variant of Diffusion Policy has an inference time of around 1.5 seconds per forward pass. Thus, we choose as our baseline method the faster and more realistic DDiM variant of Diffusion Policy, which uses 15 steps for policy inference.
We maintain the same policy inputs, outputs, and architecture as in the simulation experiment setup, except the observation size of the plug insertion task: for this task, we use image size of 256x256. At a given timestep $t$, our policy receives: an over-the-shoulder image, a camera-in-hand image, current end-effector pose and current gripper width. Our policy outputs an action sequence of length 16, where each step is a 10D vector of goal end effector position (3D), goal effector rotation (6-D) and goal gripper position (1D).
@@
For the mobile manipulation microwave task, we use a Kinova Gen3 7-DoF arm mounted on a holonomic mobile base. Inference was completed on the same laptop with a 3070 Ti GPU and transmitted to the robot over a router connection. 84x84 resolution image observations were taken from a wrist camera as well as a base mounted camera. Low dimensional observations include 3D base pose, 3D arm position, 4D arm rotation quaternion, and 1D for gripper width. Individual actions have dimension 13 (3 for the base pose, 3 for the arm position, 6D arm rotation, 1D gripper) and are chunked together temporally.
Metrics: For each trained policy, we report average success rates on one policy checkpoint (selected using validation mean-squared error). We also report inference time on the 3070 Ti GPU. For success rates, we average over 10 trials for the first and third tasks while we average over 20 trials for the second task. Starting positions were randomized between trials for the first two tasks and were static for the third task.
Results: Table IV shows how the baseline DDiM-variant of Diffusion Policy achieves similar average success rates as our method on the Rubbish Clean Up and Plug Insertion task. However, our method has much lower inference time ($\sim$9x lower latency). Our method maintains its inference speed advantage in the Microwave task and performs slightly worse than DDiM. More discussion about the mobile task in particular is present in Limitations see Sec. V.
As Consistency Policy requires 15x fewer forward passes than DDiM, one might expect the full inference speed-up to also be around 15x instead of around 9x. This discrepancy exists because of overhead costs such as the observation encoder. We discuss inference speeds further in the Appendix.
Figure 7: Microwave. This task involves: (1) navigating to and opening a microwave, (2) navigating to and picking up a bag of broccoli, (3) placing the bag inside the microwave, (4) closing the microwave door, and (5) starting the microwave.
Trash Clean Up Plug Insertion Microwave
Success Inference Success Inference Success Rate
@@
DDiM 0.8 $\pm$ .13 192 0.6 $\pm$ 0.11 198 0.5 $\pm$ 0.16
CP (ours) 0.8 $\pm$ .13 21 0.7 ±$\pm 0.10$ 22 0.4 $\pm$ 0.15
Table IV: Real World Experiment Results for Multiple Tasks – Results presented for Consistency Policy and the DDiM variant of Diffusion Policy. Success rates and standard errors are presented for each task. Inference speeds are also presented for the first two tasks.
@@ V Limitations
Our mobile manipulation results, where Consistency Policy slightly underperforms Diffusion Policy on accuracy, showcases the varied strengths and weaknesses of our method.
Note first that we were unable to train Consistency Policy on the Microwave task until it was no longer improving due to external time constraints. Note as well that Consistency Policy empirically needs more time in training to reach the same level of performance as Diffusion Policy, especially when taking into account time taken to train the teacher policy. This usually means both more epochs as well as more time per epoch since a Consistency Policy training step requires running the teacher model in addition to multiple forward passes of the student network.
This increase in training time is exacerbated by the difficulty of the task (Microwave took much longer to distill than either of the other tasks did) and is intuitive given that Consistency Policy is fitting a harder problem (a single-step policy rather than a multi-step policy over the same task). It is important to consider how these trade-offs interact with your unique use case, rather than comparing these tools in a vacuum.
@@ VI Conclusion
We present the use of consistency-based training objectives for training high performing, low-latency visuomotor robot policies. In our evaluation suite of 9 tasks in simulation and in the real world, we demonstrate that consistency policies yield a dramatic increase in inference speeds compared to previous diffusion policy methods without sacrificing success rates. Through our ablations, we also highlight key design decisions that led to visuomotor policies with the highest success rates. These include: the choice of consistency objective, lowering the initial sample variance, the use of dropout, and the choice of preset chaining steps. In future work, we hope to explore deeper explanations behind certain experimental results, such as chaining dynamics and the role of dropout along the $s\rightarrow 0$ generations, and also port Consistency Policy to robots that can better utilize fast inference speeds, including legged, winged, and other mobile platforms.
@@ Appendix B Low Variance Initial Samples
Testing EDM on unconditional image generation suggests that this low-variance sample might have some specific benefit for the robotics domain. Figures 9 and 9 are generations from the same EDM model (trained on CIFAR-10, a $32$x32323232 px image dataset) using low initial variance sampling and high initial variance sampling.
@@ Appendix C Real World Experiment Setup
We perform the stationary arm real world experiments (Trash Pick Up and Plug Insertion) using a Franka Panda robot. We mount a wrist (Zed Mini) camera and an over-the-should camera (Zed Mini) to obtain two images at each timestep. We use an Meta Quest 2 VR setup to collect 180 demonstrations for our task. The Franka robot accepts control commands at 1kHz, and receives new commands from either the VR controller or trained policy at 15 Hz. Note that Consistency Policy generates an end effector action sequence output in around 21ms, and we supply each waypoint to the robot at 15Hz. When the policy performs inference, we send a control command to the robot to maintain its current end effector pose.